Enter the quadratic equation in standard form. Use the caret (^) to enter exponents. For example, enter x2 as x^2.
x + 9 = 4(x - 1)2

Answer by Edwin McCravy(20055)

       x + 9 = 4(x - 1)2

       x + 9 = 4(x - 1)^2

       x + 9 = 4(x - 1)(x - 1)

       x + 9 = 4(x^2 - x - x + 1)

         x + 9 = 4(x^2 - 2x + 1)

         x + 9 = 4x^2 - 8x + 4

-4x^2 + 9x + 5 = 0 

 4x^2 - 9x - 5 = 0 

Edwin